OpenStudy (anonymous):

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions. s squared- 5s-6=0

OpenStudy (***[isuru]***):

let's c.. f(s) = s^2 -5s -6 when f(s) = 0 Delta S = (-5)^2 - (4 * [-6] * 1) = 25 + 24 = 49 therefor Delta S > 0 which means f(S) has two different rational solutions when f(s) = 0

OpenStudy (anonymous):

thank you. What about t^2+12t+36=0

OpenStudy (anonymous):

v^2-8v-3=0 and w^2+3w+4=0

OpenStudy (***[isuru]***):

u r welcome for ur second part let's take f(t) , such that f(t) = t^2 +12t + 36 when f(t) = 0 delta t = 12^2 - [4 * 36 ] = 144 - 144 = 0 when discriminant is equal to 0 the quadratic expression will have exactly one rational root which means f(t) is a perfect square it's actually f(t) = (t + 6)^2

OpenStudy (anonymous):

v^2-8v-3=0

OpenStudy (anonymous):

w^2+3w+4=0

OpenStudy (***[isuru]***):

u seem to have a flow of endless questions bro, let f(v) = v^2 -8v -3 delta v = (-8)^2 - [ 4 * (-3) ] = 64 + 12 = 76 which means delta v >0 which mean f(v) has two different rational solutions when f(t) = 0

OpenStudy (anonymous):

thank you. last one, w^2+3w+4=0

OpenStudy (anonymous):

is it two irrational solutions?

OpenStudy (***[isuru]***):

let f(w) = w^2 +3w +4 delta w = (3)^2 - [ 4 * (4) ] =9 - 16 = - 7 which means delta w < 0 which mean f(w) has two different irrational solutions f(w) = 0

OpenStudy (anonymous):

thank you